OpenStudy (anonymous):

Write the equation of the line that passes through (2, 2) and (6, 3) in standard form. (2 points) Question 3 options: 1) 4x + 4y = 6 2) x + 4y = 12 3) 4x + 4y = -12 4) x - 4y = -6 more help please?

OpenStudy (bahrom7893):

y-y1 = M(x-x1) M = (y-y1)/(x-x1)

OpenStudy (bahrom7893):

M = (2-3)/(2-6) = 1/4

OpenStudy (bahrom7893):

y-2=1/4(x-2)

OpenStudy (bahrom7893):

4y-8=x-2 4y-x=6 x-4y=-6

OpenStudy (anonymous):

thank youuu so much